OpenStudy (anonymous):

Can someone explain how to differentiate y=cos^2(3x)?

geerky42 (geerky42):

\(\large \cos^2(3x)\) is same as \(\large (\cos(3x))^2\) Use chain rule for this.

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

Wait, actually, I'm still a bit confused. Wouly you also take the derivative of (3x)?

OpenStudy (anonymous):

*Would

geerky42 (geerky42):

Yes.

geerky42 (geerky42):

So is everything clear?

OpenStudy (anonymous):

I think I need to have the steps be broken down, because I keep getting the wrong answer... @geerky42

geerky42 (geerky42):

Do you know how chain rule works? \[\dfrac{d}{dx}f(g(x)) = f'(g(x))\cdot g'(x)\] For function that is made of 3 parent functions; \[\dfrac{d}{dx}f(g(h(x))) = f'(g(h(x)))\cdot g'(h(x)) \cdot h'(x)\] Basically, first you derivative "outer layer", then you do "inner", then you derivative inner of inner and so on. Is that clear?

OpenStudy (anonymous):

That's perfect! Thank you! Also I have one more question, if you don't mind.

geerky42 (geerky42):

So you can see, \(\cos^2(3x)\) can be seen as \(f(g(h(x))),\) where \(f(x) = x^2,~g(x)=\cos x,~h(x)=3x\)

geerky42 (geerky42):

Sure

OpenStudy (anonymous):

How would I take the derivative of f(x)=e^(2/x)?

geerky42 (geerky42):

Chain rule again, f(x) = e^x and g(x) = 2/x

OpenStudy (anonymous):

Oh, okay! Why wouldn't it just be as it is? Because the derivative of e^x is e^x, so I'm a bit confused.

geerky42 (geerky42):

Of course, but you have to derivative inner too. So derivative of e^(2/x) is not e^(2/x)

geerky42 (geerky42):

you would have to take derivative of inner, 2/x, then multiply it to e^(2/x). Does that make sense so far?

OpenStudy (anonymous):

Yeah

OpenStudy (anonymous):

I got the answer right! Thank you for all the help!

geerky42 (geerky42):

No problem!